Points A,B,and C are collier A is between B and C AB=4x-3, BC=7x+5,and AC=5x-16. find each value. Please help me solve this problem
Answer by tommyt3rd(5050)

BA+AC=BC
(4x-3)+(5x-16)=7x+5
9x-7x=5+19
x=12
AB=45
AC=44
BC=89